Shift report 20010824 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 24 2001 - 08:23:58 EDT

  • Next message: Flemming Videbaek: "Re: Shift report 20010823 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Aug 24 08:23:57 2001, Shift: 00:00-08:00
    Supervisor: J.I.Jordre, with: A.Makeev
    Log book#8 pages 15 - 15
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beam because of power supply problems to the yellow ring.
    2. Only pedestal runs 4356 and 4360.
    3. Even without beam the ZDC scalers reported ~8 hits/s. When Dana arrived, he found that LED was on. The LED was removed. He found also that many blue cables were not properly connected, like cables labeled "Blue Low" were on input, not output.
       Had also a problem with IOServer and RunError messages in DAQ control. This was fixed by restarting IOServer and RunServer, as well as resetting all "sbe(1-6) bins" in Rack 7. This was thanks to Problem Solving Page on the shift web page.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 24 2001 - 08:24:34 EDT